# The Complexity of Maximizing the MST-ratio

### v1

@@ 3 Maximum MST-ratio of Geometric Point Sets
with probability one, where $\beta(d)$ is a constant. Thus, the average EMST-ratio over all proper bipartitions of $P$ is
$\sum_{P=B\cup R}\frac{|EMST(B)|+|EMST(R)|}{|EMST(P)|}$ $\sim\sum_{k=1}^{n-1}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
$=\sum_{k=1}^{n-1}\frac{[k^{1-1/d}+(n-k)^{1-1/d}].\binom{n}{k}}{n^{1-1/d}.(2^{n}-2)}$
@@
with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [4, 14]), and therefore converges to $d]{2}$ as $n$ goes to infinity. ∎
Certainly, Theorem 4 implies that $3$ is an upper bound for the average EMST-ratio of point sets in the Euclidean space. How much can this upper bound be improved? There exist point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio $\approx 2.154$.
[…]
@@
Figure 5: An example of a point set with a large average EMST-ratio consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin. As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
@@
For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup\{0\}$, where $A^{\prime}$ (resp. $B^{\prime}$, $C^{\prime}$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp. $b$,$c$), for $\varepsilon$ being small enough, see Figure 5. In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random two coloring, each of the sets $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$ will contain both colors.
@@
As for the lower bound, we show that the average EMST-ratio for $n$ points is always not less than $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [10].
Theorem 7.
@@
For every set of $n$ points in the Euclidean space, the average EMST-ratio is at least $\frac{n-2}{n-1}$.
Proof.
@@
Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in an Euclidean minimum spanning tree $T$ of $P$. Let us assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$. Considering a coloring of $P$ in which $R$ is the set of red points and $B=P\setminus R$ is the set of blue points, we call an edge of $T$ colorful if the two endpoints of it have different colors. Let $i$ be the smallest index such that the edge $e_{i}$ is colorful. $EMST(R)\cup EMST(B)\cup\{e_{i}\}$ is a spanning tree of $P$ and therefore, its weight is at least $|EMST(P)|$. This implies $|EMST(R)|+|EMST(B)|\geq|EMST(P)|-w_{i}$.
@@
On the other hand, considering a random coloring of $P$ with two colors, the probability for the edge $e_{i}$ being colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent. Thus, for a fixed index $j$, the probability that in a random coloring, the shortest colorful edge of $T$ is $e_{j}$ is $(\frac{1}{2})^{j}$. Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful.
By the above arguments and since we have $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
@@
$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|MST(P)|-\frac{|MST(P)|}{n-1},$
and hence the claim. ∎
Note that Theorem 7 holds also for general graphs.
The average EMST-ratio appears to be greater than 1 in many cases, though this is not always true. The examples we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-5\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 5$ points are sufficiently far from the core such that their pairwise distances exceed their distances to the core, see Figure 6. We believe those are the only point sets with an average EMST-ratio less than $1$.
Figure 6: A point set with average EMST-ratio smaller than $1$.
Question 2.
Can the point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio less than $1$ be characterized?
@@ 5 Computational Experiments
This section presents the results of computational experiments comparing the average EMST-ratio, maximum EMST-ratio, and the EMST-ratio for bipartite coloring, all conducted on random point sets. The setup is as follows. In a single experiment, we consider $n$ points selected uniformly at random in the unit square, we compute the EMST-ratios for all $22^{n}-$ 2 proper colorings of this point set, and we take the maximum and the average over all colorings. Additionally, we compute the EMST-ratio for the bipartite coloring. For $n$ between $5$ and $20$, we compute the mean values of the functions (i.e. the average EMST-ratio, maximum EMST-ratio, and bipartite EMST-ratio) across 100 experiments. The resulting averages are displayed in Figure 7. According to Theorem 6, the average EMST-ratio (the red curve) tends to $\sqrt{2}$. Interestingly, the diagram in Figure 7 suggests that a similar behavior might hold for the maximum EMST-ratio (the green curve) as well, though approaching a different limit. In fact, the program did not find any instance of a random point set with a maximum MST-ratio greater than $2$ in our experiments.
Figure 7: The mean values of the maximum EMST-ratios (green), the bipartite EMST-ratios (blue), and the average EMST-ratios (red) of $100$ random sets of $n$ points, where $5\leq n\leq 20$.
Question 3.
@@
As for the bipartite coloring, Figure 7 clearly shows how close its EMST-ratio is to the maximum EMST-ratio in expectation. Figure 8 particularly compares the EMST-ratio for the bipartite coloring and the maximum EMST-ratio across $100000$ experiments involving $5$ to $20$ points. Moreover, the maximum EMST-ratio is never more than $1.3$ times and is rarely more than $1.1$ times the bipartite EMST-ratio across all these experiments.
Figure 8: Each point in the diagram represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the maximum EMST-ratio, respectively.
Conjecture 4.
For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $1.1$ times the bipartite EMST-ratio with probability tending to $1$ as $n$ goes to infinity.
Figure 9: In this diagram each point represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the average EMST-ratio, respectively.
Figure 9, on the other hand, presents a comparison between the EMST-ratio for the bipartite coloring and the average EMST-ratio obtained from $100000$ experiments with point sets ranging from $5$ to $20$ points. Interestingly, in all computed experiments without exception, the bipartite EMST-ratio exceeds the average EMST-ratio. This suggests that the bipartite coloring of a point set provides a better EMST-ratio than what is expected from a random coloring. Although this is not always the case for finite point sets, we believe there are very few counterexamples.
Question 4.
Can the point sets in ${{\mathbb{R}}}^{2}$ for which the bipartite EMST-ratio is smaller than the average EMST-ratio be characterized?
@@ 6 Discussion
For a point set $P$ in the plane and a partition $P=R\cup B$ of $P$ into red and blue points, the Chromatic Crossing Number is defined as the number of crossings between pairs of edges, one in $EMST(R)$ and one in $EMST(B)$. Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring. In fact, for any point set, it is easy to find colorings with no chromatic crossings. More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
Question 5.
@@
Question 6.
Is there any relation between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the EMST-ratio tend to have a high chromatic crossing number?
Conjecture 5.
Given a point set $P$ in the plane, determining the maximum chromatic crossing number of $P$ is NP-hard.
@@
Related question to finding the maximum MST-ratio is the following. In the MAX-EMST-subset problem, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$. The question can be raised also in the abstract setting. Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST length over all induced subgraphs $H\subseteq G$.
Question 7.
What can we say about the complexity of MAX-MST-subset and MAX-EMST-subset problems? Are they NP-hard?
@@
Similar questions can be asked when we color a point set with $c>2$ colors. To express the MST-ratio, we consider the combined length of the minimum spanning trees of points in each color divided by the length of the minimum spanning tree of the (uncolored) point set. Theorem 4 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space.

### v2

@@ 3 Maximum MST-ratio of Geometric Point Sets
with probability one, where $\beta(d)$ is the constant in (1), and $k_{0}$ is an appropriate constant. Since $n$ is significantly greater than $k_{0}$, and the EMST-ratio is at most $3$, the contribution of the EMST-ratios for colorings with $k<k_{0}$ red points or with $n-k<k_{0}$ blue points are negligible when computing the average EMST-ratio. Thus, the average EMST-ratio over all proper bi-partitions of $P$ is
$\sum_{P=R\cup B}\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}$ $\sim\sum_{k=k_{0}}^{n-k_{0}}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
$=\sum_{k=k_{0}}^{n-k_{0}}\frac{[k^{1-1/d}+(n-k)^{1-1/d}].\binom{n}{k}}{n^{1-1/d}.(2^{n}-2)}$
@@
with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [6, 17]), and therefore converges to $d]{2}$ as $n$ goes to infinity.
@@
Figure 5: A point set consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin. The corresponding minimum spanning tree is depicted. As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
@@
Certainly, Theorem 3.8 implies that $3$ is an upper bound for the average EMST-ratio of point sets in Euclidean space. How much can this upper bound be improved? There exist point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio $\approx 2.154$. For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A\cup B\cup C\cup\{0\}$, where $A$ (resp. $B$, $C$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp. $b$,$c$), for $\varepsilon$ being small enough, see Figure 5. In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random bi-partition, each of the sets $A$, $B$ and $C$ will contain both colors.
@@
As for the lower bound, we show that the average EMST-ratio for $n$ points is always at least $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [13].
Theorem 4.3.
@@
For every set of $n$ points in Euclidean space, the average EMST-ratio is at least $\frac{n-2}{n-1}$.
Proof 4.4.
@@
Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in a Euclidean MST, $T$, of $P$. Assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$. Considering a coloring $P=R\cup B$ of $P$ into red and blue points, let $i$ be the smallest index such that the edge $e_{i}$ is colorful. $EMST(R)\cup EMST(B)\cup\{e_{i}\}$ is a spanning tree of $P$ and therefore, its weight is at least $|EMST(P)|$. This implies $|EMST(R)|+|EMST(B)|\geq|EMST(P)|-w_{i}$.
@@
On the other hand, considering a random coloring of $P$ with two colors, the probability for the edge $e_{i}$ being colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent. Thus, for a fixed index $j$, the probability that in a random coloring, the shortest colorful edge of $T$ is $e_{j}$ is $(\frac{1}{2})^{j}$. Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful. By the above argument and since there are $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
@@
$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|EMST(P)|-\frac{|EMST(P)|}{n-1},$
and hence the claim.
Note that Theorem 4.3 holds also for general graphs.
The average EMST-ratio appears to be greater than 1 in many cases, though this is not always true. The only instances we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-6\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 6$ points are sufficiently far from the core such that their pairwise distances are not less than their distances to the core. We believe those are the only point sets with an average EMST-ratio of less than $1$.
As an example, we show that the average EMST-ratio for the point set in Figure 6 is less than $1$. The point set consists of $n-5$ points in a small circle through the origin whose radius is $\varepsilon\in{{o}}(\frac{1}{n2^{n}})$ (the core), along with $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin. In this example, clearly $|EMST(P)|\geq 5-{{o}}(n\varepsilon)$. In any bi-coloring of $P$ by red and blue if an outer point has the same color as one of the core points, then the edge connecting this outer point to a core point appears in one of the monochromatic EMSTs. Therefore, in most of the bi-colorings of $P$ where the core contains both colors, we have $|EMST(R)|+|EMST(B)|\leq 5+{{o}}(n\varepsilon)$. There are only $62$ many proper bi-colorings with a mono-chromatic core, and one can show that in each of them $|EMST(R)|+|EMST(B)|<5.1$. In fact, in $52$ of such colorings we even have $|EMST(R)|+|EMST(B)|<4.8$. Thus the average EMST-ratio for this point set is upper bounded by:
$\frac{(2^{n}-64)\times(5+{{o}}(n\varepsilon))+52\times(4.8)+10\times(5.1)}{(2^{n}-2)\times(5-{{o}}(n\varepsilon))}<1.$
Figure 6: A point set with average EMST-ratio smaller than $1$. The point set consists of $n-5$ points in a very small circle through the origin (the core), and $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin. The corresponding MST is depicted.
Question 4.5.
Can the point sets in ${{\mathbb{R}}}^{2}$ with an average EMST-ratio $<1$ be characterized?
@@ 5 Computational Experiments
This section presents the results of computational experiments comparing the average EMST-ratio, maximum EMST-ratio, and the EMST-ratio for a bipartite coloring, all conducted on random point sets. The setup is as follows. In a single experiment, we consider $n$ points selected uniformly at random in the unit square, compute the EMST-ratios for all $22^{n}-$ 2 proper colorings of this point set, and take the maximum and the average over all colorings. Additionally, we compute the EMST-ratio for a bipartite coloring of the sample point set, by computing an MST using Prim’s algorithm [17]. For $n$ between $5$ and $20$, we calculate the mean values of the functions (i.e. the average EMST-ratio, maximum EMST-ratio, and bipartite EMST-ratio) across $500$ experiments. The resulting averages are displayed in Figure 7. According to Theorem 4.1, the average EMST-ratio (the red curve) tends to $\sqrt{2}$. Interestingly, the diagram in Figure 7 suggests that a similar behavior might hold for the maximum EMST-ratio (the green curve) as well, though approaching a different limit. In fact, the program did not find any instance of a random point set with a maximum EMST-ratio greater than $2$ in our experiments.
Figure 7: The mean values of the maximum EMST-ratios (green), the bipartite EMST-ratios (blue), and the average EMST-ratios (red) of $500$ random sets of $n$ points, where $5\leq n\leq 20$.
Question 5.1.
@@
As for the bipartite coloring, Figure 7 clearly shows how close its EMST-ratio is to the maximum EMST-ratio in expectation. Figure 8 particularly compares the EMST-ratio for a bipartite coloring computed using Prim’s algorithm and the maximum EMST-ratio across $10^{5}$ experiments involving $5$ to $20$ points. Moreover, the maximum EMST-ratio is never more than $1.3$ times and is rarely more than $1.1$ times the computed bipartite EMST-ratio across all these experiments.
Figure 8: Each point in the diagram represents a random point cloud, with the $x$ and $y$ coordinates corresponding to a bipartite EMST-ratio of the sample point set computed using Prim’s algorithm and the maximum EMST-ratio, respectively.
Conjecture 5.3.
For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $1.1$ times any bipartite EMST-ratio of the point set with probability tending to $1$ as $n$ goes to infinity.
Figure 9: In this diagram each point represents a random point cloud, with the $x$ and $y$ coordinates corresponding to a bipartite EMST-ratio of the sample point set computed using Prim’s algorithm and the average EMST-ratio, respectively.
Figure 9, on the other hand, presents a comparison between a bipartite EMST-ratio computed using Prim’s algorithm and the average EMST-ratio obtained from $10^{5}$ experiments with point sets ranging from $5$ to $20$ points. Interestingly, in all computed experiments without exception, the computed bipartite EMST-ratio exceeds the average EMST-ratio. This suggests that bipartite colorings of a point set provide better EMST-ratios than what is expected from a random coloring. Although this is not always the case for finite point sets, we believe there are very few counterexamples.
Question 5.4.
Can the point sets in ${{\mathbb{R}}}^{2}$ for which a bipartite EMST-ratio of them is smaller than the average EMST-ratio be characterized?
@@ 6 Discussion
For a point set $P$ in the plane and a partition $P=R\cup B$ of $P$ into red and blue points, the Chromatic Crossing Number is defined as the number of crossings between pairs of edges, one in $EMST(R)$ and one in $EMST(B)$. Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring appears. In fact, for any point set, it is easy to find colorings with no crossings between the two MSTs. More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
Figure 10: An example of a point set and a bi-partition with chromatic crossing number $4$.
Question 6.1.
@@
Question 6.2.
Are there relations between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the chromatic crossing number tend to have a relatively high EMST-ratio?
Conjecture 6.3.
Finding the maximum chromatic crossing number of a point set in the plane is NP-hard.
@@
Related question to finding the maximum MST-ratio is the following. In the MAX-EMST-subset problem, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$. The question can be raised also in the abstract setting. Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST weight over all induced subgraphs $H\subseteq G$.
Question 6.4.
Are the problems MAX-MST-subset and MAX-EMST-subset NP-hard?
@@
Similar questions can be asked for point sets with $c>2$ colors. To express the MST-ratio, we consider the total length of minimum spanning trees of points in each color divided by the length of a minimum spanning tree of the (uncolored) point set. Theorem 3.8 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space. For points in Euclidean space, this bound can be improved for $c\geq 4$. In particular, as $\rho_{d}\geq 0.577$ for every $d$, the maximum EMST-ratio is upper bounded by $\frac{c}{0.577}$. It remains unclear how the results in [9] on lattices can be generalized to more than $2$ colors. Additionally, exploring the average EMST-ratio for $c$ colors is an interesting direction to pursue.